OpenStudy (harleygirl18):

Help please no guessing please explain how to solve Will medal. Julie earned simple interest on $8250 with an annual interest rate of 3.9% for 7 years. How much did she earn in interest? Round your answer to the nearest cent. simple interest: I = P • r • t Answer_______

OpenStudy (harleygirl18):

@Zeronknight @zaynab123 @kittiwitti1 @LoveIt @ganeshie8 @Preetha

OpenStudy (harleygirl18):

@surjithayer

OpenStudy (calculusxy):

Use the formula of \(\text{ Simple Interest = Principal x Rate x Time}\)

OpenStudy (simplymarie_x):

t = time r = rate p = principal so I = 8250 * 3.9 * 7.

OpenStudy (calculusxy):

Principal = 8250 (what you're starting with) Interest = 0.039 (the decimal form of 3.9%) Time = 7 years

OpenStudy (anonymous):

P=$ 8250 r=3.9%\[=\frac{ 3.9 }{ 100 }=\frac{ 39 }{ 1000 }\] t= 7 years plug the values and calculate interest.

OpenStudy (calculusxy):

Thus you have \(\large I = $8250 \times 0.039 \times 7\)

OpenStudy (simplymarie_x):

^^

OpenStudy (harleygirl18):

so the answer 2252.25

OpenStudy (calculusxy):

Correct!

OpenStudy (harleygirl18):

thanks can you help with two more

OpenStudy (calculusxy):

I'll try

OpenStudy (harleygirl18):

With a principal investment of $25,000, which account will have the greatest value after 6.5 years? simple interest: I = P • r • t interest compounded annually is A = P (1 + r)t interest compounded quarterly: A = P (1 + r/4 )4t A. 2.6% with interest compounded quarterly B. 3.2% in a simple interest account C. 2.9% with interest compounded annually D. 2.3% in a simple interest account

OpenStudy (raffle_snaffle):

well what do you think?

OpenStudy (calculusxy):

use each of the answer choices' information. a) 2.6% with interest compounded quarterly A = P (1 + r/4 )4t A = \(25000(1 + 0.026 \div 4)^{4(6.5)}\) Solve this. use the same approach by following the equation and replacing the numbers accordingly.

OpenStudy (harleygirl18):

Wow thats a lot more easier than I thought,Just one more question and I'll be out of your hair.

OpenStudy (calculusxy):

ok

OpenStudy (harleygirl18):

Skylar plans to use $3400 to open a savings account with an annual interest rate of 1.15%. How much more interest will he earn over 13 years if he chooses a compound interest account that compounds interest quarterly instead annually? Round your answer to the nearest cent. interest compounded annually: A = P (1 + r)t interest compounded quarterly: A = P (1 +r/4)4t

OpenStudy (calculusxy):

Just substitute the numbers into their respective places in the formula. Principal = $3400 Interest Rate (in decimals) = 0.015 Time = 13 years

OpenStudy (calculusxy):

Then find which one will be the greater amount.

OpenStudy (raffle_snaffle):

Because there is 4 quarters in one year that should give you an idea which one will be larger